# CBQ: Cross-Block Quantization for Large Language Models

### v4

@@ 3.2 Coarse-to-fine pre-processing
Table 1: Evaluation on multiple zero-shot datasets with the accuracy $\uparrow$ metric, where the Mutual dataset is evaluated with the Mean Reciprocal Rank/Recall@$1$/Recall@$2$ metrics. CBQ represents that the experiments conducted with 2-bit weight-only quantization did not involve fully quantizing the model but only the FC2 layera of the first and the last transformer blocks are converted to 4-bit precision.
@@
Table 2: Evaluation quantization on generation datasets with the perplexity (PPL) $\downarrow$ metric, where OmniQ represents OmniQuant.
To address the challenges in achieving optimal quantization parameters before cross-block reconstruction, we have conducted a thorough analysis of outliers in both weights and activations. Based on this analysis, we have developed an effective and unified design approach to handle extreme weight and activation outliers.
Figure 4 illustrates the block-by-block visualization of the weight distribution of the last fully connected layer, as well as the channel-by-channel visualization of the activation distribution within one block. It is observed that outliers in both weights and activations contribute to larger reconstruction errors between the quantized and full-precision models. To address this issue, we analyze the distribution of outliers and propose a coarse-to-fine pre-processing strategy to suppress outliers and reduce the difficulty of reconstruction.
Coarse-grained detection.
The comprehensive algorithm of the outlier detection is illustrated in Algorithm 1, which is divided into two stages. In the first stage, we perform coarse-grained detection by calculating the lower and upper quartile values ($Q_{1}$ and $Q_{3}$) and the interquartile range ($IQR$) in the numerical distribution (either activations or weights). Based on these calculations, we obtain a coarse outlier set $O=\{x|x>T,x\in X\}$, where $T=Q_{3}+\lambda_{1}IQR$ and $\lambda_{1}$ is set to 1.5. This stage greatly reduces the search space for outlier detection.
Fine-grained detection.
Removing outliers in weights has minimal impact on performance, whereas outliers in activations, particularly in specific channels, can greatly affect performance if directly removed. Consequently, our approach involves truncating weight outliers and scaling outliers in activations based on the detected outliers in both weights and activations. This is different from existing outlier pre-processing methods (Xiao et al., 2022; Wei et al., 2023) that focus solely on activations. Figure 4 provides visual evidence of weight outliers being truncated within the outlier group. Additionally, the figure demonstrates the presence of outliers in specific channels. Scaling these specific channels results in a significant concentration of the overall distribution.
$s_{i}=\sqrt{\textrm{Max}(|X_{i}|)/\textrm{Max}(O^{*})}.$ (6)
This scaling factor is then applied to update weights and activations following prior work (Wei et al., 2023) to counteract destabilizing fluctuations from remaining outliers.
3.3 LoRA-Rounding for weight quantization
The performance degradation raised by weight quantization is non-negligible when compressing the model into low bit-width. The quantization error of weights comes from the rounding error (Nagel et al., 2020) and clipping error (Tu et al., 2021), but previous PTQ methods for LLMs only focus on the optimization of the latter one and do not take the rounding error into consideration. AdaRound (Nagel et al., 2020) introduces to learn a better weight-rounding matrix for post-training quantization that adapts to the data and the task loss. As shown in Eq. 7, we can obtain the weight-rounding matrix $\Delta_{W}\in\mathbb{R}^{d\times k}$ with a learnable matrix $V\in\mathbb{R}^{d\times k}$ with a rectified sigmoid function:
However, the technique of AdaRound is hard to apply directly to LLMs with billion-level parameters, which could lead to much computational overhead and significantly reduce the efficiency of post-training quantization. To further rectify weight quantization errors arising from rounding operations and prevent their accumulation in adjacent blocks, we introduce an efficient adaptive rounding method for LLMs, called LoRA-Rounding. Specifically, inspired by LoRA (Hu et al., 2021), we employ low-rank adaptive learning on the weight-rounding matrices, decomposing $V$ with much smaller low-rank matrices, and only optimize them in post-training quantization, the decomposition is defined as:
where the rank $r<<\textrm{min}(d,k)$. Following the setting of LoRA (Hu et al., 2021), we utilize a random Gaussian initialization for $A_{1}$ and zero for $A_{2}$, thus $\Delta_{W}$ is set to zero at the beginning of post-training quantization. During training, each element of $V$ is encouraged into 0 or 1 with a regularizer loss:
$\mathcal{L}_{round}=\sum_{i,j}1-|2\Delta_{W}{(i,j)}-1|^{\beta},$ (9)
where $\beta$ is a annealing factor. Following AdaRound (Nagel et al., 2020), $\beta$ is set to higher in the initial phase and set to lower in the later phase of the optimization to encourage it to converge to 0 or 1. We also conduct $\Delta_{W}=\left\lfloor\Delta_{W}\right\rceil$ in the later phase of the optimization to force each element into $\{0,1\}$ exactly.
Compared with vanilla AdaRound for LLMs. The proposed LoRA-Rounding reduces the number of learnable parameters from $d\times k$ to $(d+k)\times r$ and accelerates the optimization process significantly, we conduct ablation experiments in the next section.
Compared with QLoRA (Dettmers et al., 2023). QLoRA is an efficient parameter finetuning method for quantized LLMs, freezes the quantized weight and optimizes the float low-rank matrices, which is much similar to the original LoRA but takes lower memory. However, LoRA-Rounding only optimizes the rounding error of weight quantization, which is a much different method that is designed for post-training quantization with ultra-low cost.
3.4 Overall loss
In summary, by utilizing a low-rank decomposition of the weight-compensated matrix, we effectively reduce the number of learnable parameters $S_{W},S_{X},A_{1},A_{2}$ while still achieving accurate correction of weight quantization errors within each block. The total loss for optimizing the $i^{th}$ block to the $k^{th}$ block within a sliding window is formulated as
$\mathcal{L}_{total}=\mathcal{L}_{rec}+\gamma\mathcal{L}_{round},$ (10)
where the $\gamma$ is the hyper-parameter to balance the reconstruction error and round error, which is set to 0.01 in our experiments. During optimization, We slide the window to the last block with an interval and update all the quantization parameters within the window, ensuring that the cumulative error in the quantized transformer block is minimized and achieves optimal performance.
4 Experiments
In this section, we conduct extensive experiments to validate the effectiveness and superiority of the proposed CBQ.
4.1 Setup
Models and datasets.
We conduct experiments on large language models with different sizes, including OPT (Zhang et al., 2022) and LLAMA (Touvron et al., ) families. We validate our quantization scheme on various datasets which are divided into two categories. One is reported by the perplexity metric of language generation experiments on C4 (Raffel et al., 2020) and WikiText2 (Merity et al., 2016). The other is reported by the accuracy metric of zero-shot language tasks on PIQA (Bisk et al., 2020a), HellaSwag (Clark et al., 2018), ARC (Clark et al., 2018), Mutual (Cui et al., 2020) and Ehics (Hendrycks et al., 2020a).
@@ 4.1 Setup
To thoroughly evaluate performance, we test extensive quantization schemes including weight-only quantization down to W4A16 and W2A16, as well as joint weight-activation quantization for ultra-low bitwidths like W4A8, and W4A4. This extensive assessment across varying bitwidths provides a robust analysis of our proposed method. Also, with full consideration of the convenience of actual deployment of the quantized model, except for the settings with activation values below 4 bits that use per-token quantization, we perform per-tensor quantization on activations.
@@
For weight-only quantization settings, we selected GPTQ (Frantar et al., 2022b) as the baseline quantization method in our experiments. This represents the most prevalent technique for W4A16 quantization of language models, respectively. Furthermore, we compare our CBQ with OmniQuant (Shao et al., 2023) and QLLM (Liu et al., 2023), which is the state-of-the-art method based on block reconstruction. We include a comparison of our CBQ method with the groupwise quantization method RPTQ (Yuan et al., 2023), which is widely employed in the W4A8 setting.
@@
Following the setting of previous work (Frantar et al., 2022b), our calibration dataset comprises 128 randomly selected 2048-token segments from C4 to ensure standardized benchmarking. To balance quantization performance and training speed, we utilize sliding windows containing two blocks with 3 epochs per window. For the LoRA-Rounding technique, we set the rank $r$ to 5. The optimization process involves adjusting the learnable quantization step sizes ($S_{X}$ and $S_{W}$) and the weight-rounding matrix ($\delta_{W}$) with learning rates of $61e-$ 6, $71e-$ 7, and $41e-$ 4, respectively. To manage the learning rate, we utilize the CosineAnnealingLR scheduler. We quantize all models using a mini-batch size of 1 on a single GPU. This configuration allows efficient cross-block dependency modeling while sufficiently propagating information across windows.
@@ 4.2 Experimental results
Results on multiple zero-shot benchmarks using accuracy as the evaluation metric demonstrate CBQ’s capabilities on large models including OPT (30B, 66B) and LLAMA (30B, 65B) (as shown in Table 1). Across almost all public datasets, CBQ outperforms existing quantization methods by over 2% and reduces the accuracy gap with the full precision model to within 1% under the W4A16, W2A16 and W4A8 quantization settings. This demonstrates stronger zero-shot capability. Moreover, unlike current techniques, CBQ uniquely achieves ultra-low quantization down to W4A4 bitwidths while maintaining a higher performance than the state-of-the-arts. The consistent gains verify the generalization of CBQ’s innovations across models and datasets.
@@
#Num of blocks w/ Overlap C4$\downarrow$ Wiki $\downarrow$
1 14.57 11.98
2 14.23 11.35
13.29 10.63
4 14.32 11.45
13.27 10.60
Table 3: Ablation of the cross-block dependency (CBD).
@@
w/ Smoothquant (Xiao et al., 2022) 33.21 25.26
@@
w/ CFP-Weight + CFP-Activation + CBQ-Recon. 13.29 10.63
Table 4: Ablation on the coarse-to-fine pre-processing. ‘CBQ-Recon.’ represents the proposed cross block reconstruction.
Comparison of block reconstruction-based methods.
We provide a comprehensive validation of the superiority of our proposed CBQ method compared to the state-of-the-art block reconstruction-based methods, OmniQuant and QLLM. To evaluate the performance improvement, we conducted experiments that encompassed multiple zero-shot tasks and generation tasks. The results, as presented in Table 1 and Table 2, consistently demonstrate the significant performance gains achieved by our CBQ method.
4.3 Ablation Study
To analyze the contribution of each component in our proposed CBQ method, we performed ablation experiments on the LLAMA-7B model.
Cross-block dependency.
To analyze the impact of our proposed cross-block dependency (CBD) method, we performed ablation experiments in Table 3. Results demonstrate performance gains as the number of blocks jointly processed per sliding window increases, validating CBD’s ability to model inter-block dependencies. Furthermore, utilizing overlapping blocks between adjacent sliding windows supplements cross-window relational representation. This redundancy helps capture nuanced block interactions and enables additional accuracy improvements. Overall, these ablation studies highlight the benefits of CBD for progressively accumulating and propagating cross-block knowledge during CBQ optimization.
LoRA-Rounding.
As shown in Table 5, the incorporation of the LoRA-Rounding technique into the learning process effectively mitigates the rounding errors that typically arise from weight quantization. This improvement leads to a significant enhancement in quantization accuracy. In comparison to the conventional ’Rounding’ technique, our LoRA-Rounding approach utilizes low-rank decomposition to reduce the number of learnable parameters. This reduction not only reduces GPU memory consumption during training but also improves training speed. Despite the additional training costs, LoRA-Rounding significantly outperforms ’Rounding’ in terms of performance.
@@ 4.3 Ablation Study
Table 5: Ablation of the LoRA-Rounding.
@@
Ablation of the proposed coarse-to-fine pre-processing verifies performance improvements from suppressing outlier values before quantization, as shown in Table 4. Detecting and dynamic scaling extreme activation outliers reduces distortions, while additional truncating of weight outliers further stabilizes training. This tailored preprocessing stabilizes the data distribution for robust optimization at ultra-low bitwidths. It is important to note that this two-pronged pre-processing method is specifically designed to address the abundance of outliers in language models, which are not adequately addressed by the existing pre-processing technique of Smoothquant (Xiao et al., 2022). Furthermore, we conduct a reconstruction optimization process, referred to as ’CBQ-Recon.’, based on the pre-processed weights and activations. This approach, compared to the optimization based on the processing of Smoothquant, yields better performance.
Loss functions.
KL loss L2 loss Homologous C4 Wiki
13.82 11.13
13.84 11.12
13.55 10.88
13.29 10.63
Table 6: Ablation study on block-wise reconstruction loss functions. ‘homologous’ represents the homologous reconstruction.
To determine optimal loss formulations, we evaluate reconstruction errors using L2 alone, KLD alone, and a combination of them in Table 6. Ablation results demonstrate superior performance from KLD over L2, with the combined loss achieving further gains. This highlights the benefits of KLD for matching full-precision block distributions during CBQ optimization. Fusing both losses enables jointly minimizing absolute and divergence-based errors to improve overall block-wise reconstruction. Our analysis verifies the advantage of blended L2+KLD loss for robustly optimizing blocks as interdependent components. Moreover, the introduced homologous reconstruction scheme further reduces the reconstruction difficulty and improves the quantization performance.
Comparison of quantization efficiency.
LLAMA 7B 13B 30B 65B
OmniQuant 1.6h 3.3h 7.3h 14.4h
QLLM* 1.05 1.79 - 9.05
CBQ 0.9h 1.45 2.1h 4.3h
Table 7: Comparison of training (GPU Hours) time of our CBQ with OmniQuant. * represents that the training time of QLLM is implemented on LLAMA2 models.
We evaluate the quantization efficiency of our CBQ method and compare it to OmniQuant and QLLM, both of which are reconstruction-based PTQ methods. The GPU training hours for both methods are shown in Table 7. The results demonstrate that the training cost of CBQ can be faster than OmniQuant, particularly for larger models. This indicates that our CBQ method offers an advantage in terms of training efficiency.
@@ 5 Conclusion
In this study, we introduce CBQ, a unified post-training quantization method specifically designed for large language models. By jointly optimizing weight and activation quantization parameters through block-wise reconstruction, CBQ achieves efficient quantization without hand-craft searching. The introduced cross-block dependency scheme, low-rank decomposition scheme, and outlier detection strategy further improve the quantization performance. Experimental results demonstrate the state-of-the-art performance of CBQ, particularly in ultra-low bit settings like W4A4 and W2A16, and better maintain the capability of the full-precision model in W4A8 and W4A16 settings. The proposed CBQ method contributes to advancing PTQ techniques for LLMs, paving the way for efficient and effective deployment of large language models in resource-constrained devices.
6 Acknowledgement
@@
Appendix A Appendix
A.1 Overview
@@
GPTQ (Frantar et al., 2022b) ✓/ ✗
RPTQ (Yuan et al., 2023) ✓/ ✓
OS+ (Wei et al., 2023) ✓/ ✓
SmoothQuant (Xiao et al., 2022) ✓/ ✓
OmniQuant (Shao et al., 2023) ✓/ ✓
QLLM (Liu et al., 2023) ✓/ ✓
@@
Table 8: Comparison of different quantization methods for LLMs.
@@
A.2 Evaluation quantization for a series of OPT models

### v5

@@ 5 Experiments
5.1 Setup
Models and datasets.
We conduct experiments on large language models with different sizes, including OPT (Zhang et al. (2022)) and LLAMA (Touvron et al. ) families. We validate our quantization scheme on various datasets which are divided into two categories. One is reported by the perplexity metric of language generation experiments on C4 (Raffel et al. (2020)) and WikiText2 (Merity et al. (2016)). The other is reported by the accuracy metric of zero-shot language tasks (Gao et al. (2021)) on PIQA (Bisk et al. (2020a)), HellaSwag (Clark et al. (2018)), ARC (Clark et al. (2018)), Mutual (Cui et al. (2020)) and Ehics (Hendrycks et al. (2020a)).
Table 2: Evaluation quantization on generation datasets with the perplexity (PPL) $\downarrow$ metric, where ‘OmniQ’ represents OmniQuant.
@@ 5.1 Setup
To thoroughly evaluate performance, we test extensive quantization schemes including weight-only quantization down to W4A16 and W2A16, as well as joint weight-activation quantization for ultra-low bitwidths like W4A8, and W4A4. This extensive assessment across varying bitwidths provides a robust analysis of our proposed method. Also, In alignment with prior research (Shao et al. (2023); Liu et al. (2023a); Frantar et al. (2022c)), we use per-channel weight quantization and per-token activation quantization.
@@
For weight-only quantization settings, we selected GPTQ (Frantar et al. (2022b)) as the baseline quantization method in our experiments. This represents the most prevalent technique for W4A16 quantization of language models. Furthermore, we compare our CBQ with OmniQuant (Shao et al. (2023)) and QLLM (Liu et al. (2023a)), which is the state-of-the-art method based on block reconstruction. We include a comparison of our CBQ method with the groupwise quantization method RPTQ (Yuan et al. (2023)), which is widely employed in the W4A8 setting.
@@
Following the setting of previous work (Frantar et al. (2022b); Liu et al. (2023b); Yao et al. (2024); Yuan et al. (2023)), our calibration dataset comprises 128 randomly selected 2048-token segments from C4 to ensure standardized benchmarking. To balance quantization performance and training speed, we utilize sliding windows containing two blocks with 3 epochs per window. For the LoRA-Rounding technique, we set the rank $r$ to 5. The optimization process involves adjusting the learnable quantization step sizes ($S_{X}$ and $S_{W}$) and the weight-rounding matrix ($\delta_{W}$) with learning rates of $41e-$ 4, $31e-$ 3, and $41e-$ 4, respectively. To manage the learning rate, we utilize the CosineAnnealingLR scheduler. We quantize all models using a mini-batch size of 1 on a single GPU. This configuration allows efficient cross-block dependency modeling while sufficiently propagating information across windows.
@@ 5.2 Experimental results
Results on multiple zero-shot benchmarks using accuracy as the evaluation metric demonstrate CBQ’s capabilities on LLMs including OPT (30B, 66B) and LLAMA (30B, 65B) (as shown in Table 1). Across almost all datasets, CBQ outperforms existing quantization methods by over 2% and reduces the accuracy gap with the full precision model to within 1% under the W4A16, W2A16 and W4A8 quantization settings. This demonstrates stronger zero-shot capability. Moreover, unlike current techniques, CBQ uniquely achieves ultra-low quantization down to W4A4 while maintaining a higher performance than the state-of-the-arts. The consistent gains verify the generalization of CBQ’s innovations across models and datasets.
@@
5.3 Ablation Study
To analyze the contribution of each component in our proposed CBQ method, we performed ablation experiments on the LLAMA-7B model under W4A4.
Table 3: Ablation studies on the proposed CBD, CFP and LoRA-Rounding.
(a) Ablation of the CFP
@@ 5.3 Ablation Study
w/ OMSE (Choukroun et al. (2019)) 76.43 47.81
w/ Percentile (Zhou et al. (2017)) 71.62 45.86
w/ OS (Wei et al. (2022b)) 41.57 26.36
w/ Smoothquant (Xiao et al. (2022)) 33.21 25.26
@@
w/ CFP-Weight+Act + CBQ-Recon. 13.29 10.63
(b) Ablation of the LoRA-Rounding
@@
(c) Ablation on the CBD
[…]
1 0 14.57 11.98 17.2
2 0 14.23 11.35 21
1 13.29 10.63 21
4 0 14.32 11.45 39
1 13.27 10.60 39
2 12.56 9.56 39
3 12.32 9.45 39
Cross-block dependency.
To analyze the impact of our proposed CBD method, we performed ablation experiments in Table 3(c)c. Results demonstrate performance gains as the number of blocks jointly processed per sliding window increases, validating CBD’s ability to model inter-block dependencies. Furthermore, utilizing overlapping blocks between adjacent sliding windows supplements cross-window relational representation. This redundancy helps capture nuanced block interactions and enables additional accuracy improvements. Overall, these ablation studies highlight the benefits of CBD for progressively accumulating and propagating cross-block knowledge during CBQ optimization. For additional experimental results, please refer to Appendix  D and  E.
LoRA-Rounding.
As shown in Table 3(c), ’w/ Rounding’ indicates a modification in the training strategy for the compensation matrix, compared to the traditional ’w/ Adarounding’ approach. This change leads to a significant improvement in accuracy. Overall, LoRA-Rounding leverages low-rank decomposition to reduce the number of learnable parameters and adjusts the training strategy, which not only decreases GPU memory consumption but also enhances training speed.
@@
As shown in Table 3(a) and Table 10 in Appendix F, CFP demonstrates advantages in both weight-activation quantization and weight-only quantization. Furthermore, we conduct a reconstruction optimization process, referred to as ’CBQ-Recon.’, based on the pre-processed weights and activations. This two-pronged pre-processing effectively reduces outliers which are not adequately handled by existing preprocessing techniques like OS (Wei et al. (2022b)), Smoothquant (Xiao et al. (2022)) etc.
@@ 6 Conclusion
In this work, we conduct a detailed analysis of error sources in LLMs under low-bit quantization and identify the critical role of intra-layer and inter-layer dependencies. To address these challenges, we propose CBQ, a novel method that employs a cross-block reconstruction strategy alongside Lora-Rounding compensation matrices. This approach effectively establishes long-range inter-layer dependencies while capturing comprehensive intra-layer dependencies, surpassing traditional layer-wise and block-wise reconstruction techniques. Additionally, we introduce CFP, a technique designed to simultaneously detect and manage outliers in both weights and activations. Our experimental results demonstrate that CBQ significantly outperforms existing PTQ, achieving substantial improvements in ultra-low bit precision across a variety of tasks, while also offering enhanced computational efficiency by reducing training resource demands.
Acknowledgement
@@
Appendix A Overview
Table 4: Comparison of different quantization methods for LLMs.
@@ Appendix A Overview
GPTQ (Frantar et al. (2022b)) ✓/ ✗
RPTQ (Yuan et al. (2023)) ✓/ ✓
OS+ (Wei et al. (2023)) ✓/ ✓
SmoothQuant (Xiao et al. (2022)) ✓/ ✓
OmniQuant (Shao et al. (2023)) ✓/ ✓
QLLM (Liu et al. (2023a)) ✓/ ✓
@@
Appendix B Ablation on the loss functions
Table 5: Ablation study on block-wise reconstruction loss functions.
KL loss L2 loss C4 Wiki
13.82 11.13
13.84 11.12
13.29 10.63
To determine optimal loss formulations, we evaluate reconstruction errors using L2 alone, KLD alone, and a combination of them in Table 5. Ablation results demonstrate superior performance from KLD over L2, with the combined loss achieving further gains. This highlights the benefits of KLD for matching full-precision block distributions during CBQ optimization. Fusing both losses enables jointly minimizing absolute and divergence-based errors to improve overall block-wise reconstruction. Our analysis verifies the advantage of blended L2 + KLD loss for robustly optimizing blocks as interdependent components.
Appendix C The capability of CBQ on the LLAMA2-7B
To demonstrate the effectiveness of CBQ, we evaluate CBQ on the LLAMA2-7B model across various datasets and observed that it delivers excellent results.
Table 6: Evaluation on multiple zero-shot datasets and generation datasets on the LLAMA2-7B
#Bits Methods PIQA HellaSwag ARC-C ARC-E Mutual Ethics c4 ↓ Wiki ↓
FP - 76.93 72.95 40.69 53.21 70.92/51.12/75.84 52.63 6.97 5.47
W4A16 OmniQuant 77.14 71.86 40.18 53.70 70.00/50.46/74.74 53.10 7.12 5.58
CBQ 77.34 72.23 40.22 53.66 70.49/50.90/74.83 53.13 7.05 5.52
W3A16 OmniQuant 75.91 70.95 38.71 51.89 69.12/48.33/72.65 52.58 7.75 6.03
CBQ 76.25 71.34 39.21 52.36 69.35/49.02/73.15 52.67 7.56 5.89
W2A16 OmniQuant 68.71 53.43 30.88 39.81 65.12/42.21/69.18 50.54 12.72 9.62
CBQ 71.59 60.28 32.93 45.74 66.22/44.35/69.63 57.22 11.30 8.01
W4A4 QLLM 67.68 58.45 30.89 44.4 - - 13.26 11.75
OmniQuant 65.94 53.53 30.80 43.94 64.83/41.87/68.84 47.29 18.39 14.61
CBQ 68.25 57.34 31.56 46.23 64.89/41.87/68.74 47.59 12.56 11.32
W4A8 CBQ 76.85 72.06 40.16 53.34 70.23/50.12/74.89 52.56 7.12 5.72
W6A6 OmniQuant 76.82 72.13 39.33 53.36 69.57/49.67/73.62 52.62 7.48 5.87
CBQ 77.58 72.14 40.27 53.87 70.16/50.22/74.83 53.02 7.24 5.67
Appendix D The potential for scaling with CBD
To further investigate the scaling potential of cross-block dependency (CBD), we conducted additional experiments to explore whether increasing its scale could lead to further performance improvements.
Table 7: The scaling capability of CBQ on the LLAMA-7B under W4A4
[…]
1 0 14.57 11.98
2 0 14.23 11.35
1 13.29 10.63
4 0 14.32 11.45
1 13.27 10.60
2 12.56 9.56
3 12.32 9.45
8 0 13.56 10.78
4 11.91 9.01
7 11.86 8.96
Table 8: The capability of CBD on the LLAMA2-7B
# Num of blocks Overlap LLAMA2-7b-W2A16 LLAMA2-7b-W4A4
C4 Wiki C4 Wiki
1 0 12.34 9.12 14.28 12.33
2 0 11.89 8.76 13.85 11.96
1 11.30 8.01 12.56 11.32
4 0 11.32 8.05 12.52 11.35
1 11.12 7.95 12.15 11.01
2 11.08 7.89 11.85 10.83
3 10.92 7.82 11.5 10.62
As shown in the Table 7,8, we validate the proposed cross-block quantization in the W2A16 and the W4A4 experimental settings. Our ablation analysis in these settings underscores the robustness and versatility of CBD, showcasing its inherent simplicity to ensure seamless deployment across various quantization settings.
Appendix E Efficiency of the CBD
In order to further study the various performances with CBD, we performed the following experiments. This table below illustrates the training time, GPU memory usage, and the number of cross-block dependencies employed in the W2A16 quantization of the LLAMA-7B model.
Table 9: Ablation of the cross-block dependency (CBD) with W2A16.
#Num of blocks Overlap C4$\downarrow$ Wiki $\downarrow$ time (h) GPU memory(GB)
1 0 12.72 9.62 1.09 17.2
2 0 12.56 9.34 1.50 21
1 12.30 8.87 3.02 21
4 0 12.34 8.89 1.10 39
1 11.63 8.59 1.40 39
2 11.42 8.28 1.96 39
3 11.21 8.08 2.60 39
Our CBD considers dependencies between two blocks within a sliding window, distinguishing it from existing methods that focus solely on individual block dependencies. This unique design yields significant performance improvements while incurring additional GPU overhead. Additionally, by incorporating overlapping windows, CBD enhances cross-block dependencies without requiring additional GPU memory usage.
Appendix F The theoretical foundation of CFP
Existing outlier detection methods often assume that data follow a normal distribution, which is not always strictly applicable to real-world datasets. Our approach avoids assuming specific data distributions, providing flexibility in capturing outliers across diverse datasets. The quartile criterion is robust to outliers, as it is not heavily influenced by extreme values.
We give two commonly used methods as follows:
[…]
  • Percentile-based method: This method uses percentiles to detect outliers.
Our quartile criterion follows the existing analysis (Massart et al. (2005)), which includes the maximum value, minimum value, median, and upper and lower quartiles, to detect outliers. This approach does not require any assumptions about the distribution of the data and does not impose any restrictive requirements on the data. It simply portrays the true shape of the data, providing an objective way to identify outliers.
Table 10: The capability of CFP on the LLAMA2-7B
Method C4 $\downarrow$ Wiki $\downarrow$
w/o outlier pre-processing 1082.68 1128.33
w/ OMSE (Choukroun et al. (2019)) 76.43 47.81
w/ Percentile (Zhou et al. (2017)) 71.62 45.86
w/ OS (Wei et al. (2022b)) 41.57 26.36
w/ Smoothquant (Xiao et al. (2022)) 33.21 25.26
w/ CFP-Activation 23.48 19.75
w/ CFP-Weight + CFP-Activation 21.98 17.95
w/ OMSE + CBQ-Recon. 25.34 19.53
w/ Percentile + CBQ-Recon. 25.62 19.45
w/ OS + CBQ-Recon. 17.83 13.89
w/ Smoothquant + CBQ-Recon. 15.69 12.24
w/ CFP-Weight+Act + CBQ-Recon. 13.29 10.63
Appendix G Comparison of quantization efficiency
Table 11: Comparison of training (GPU Hours) time of our CBQ with OmniQuant.
LLAMA 7B 13B 30B 65B
OmniQuant 1.1h 2.2h 4.5h 8.9h
CBQ 0.9h 1.45 2.1h 4.3h
We evaluate the quantization efficiency of our weight-only CBQ quantization method and compare it to OmniQuant which is the representative reconstruction-based PTQ methods. The GPU training hours for both methods are shown in Table 11. The results demonstrate that the training cost of CBQ can be faster than OmniQuant, particularly for larger models. This indicates that our CBQ method offers an advantage in terms of training efficiency.
Appendix H Ablation of the rank of LoRA-Rounding
The ablation of the rank of LoRA-Rounding is in the Table 12 below. It is observed that with lower ranks (3, 4, and 5), there is a slight improvement in performance. However, as the rank increases beyond 5, the performance starts to decline. Considering the limited training resources available for LLMs, it is worth noting that a larger rank in LoRA-Rounding results in a higher number of parameters that need to be optimized. This increased parameter complexity poses a significant challenge and ultimately leads to poorer performance and results.
Table 12: Ablation of the rank of LoRA-Rounding.
Dataset Rank = 3 Rank = 4 Rank = 5 Rank = 6 Rank = 7
C4$\downarrow$ 13.4 13.35 13.29 13.46 13.98
[…]
Appendix I Evaluation quantization for a series of OPT models
To further demonstrate the performance of the proposed CBQ, we conducted additional evaluations under the OPT models as shown in Table 13.